Ethanol (C2H6O) is used as an additive in gasoline . Calculate the molality of an ethanol solution prepared by dissolving 78.6 g of ethanol in enough water to produce 225 g of solution .

Select one:

A. 11.67 m

B. 7.59 m

C. 349.33 m

D. 0.349 m

E. 0.537 m
Answer #1

Molan mass of ethanol (Catho) = (12+ 2 + 6x 1 + phase= (24+ 6+ (6) glamol = 46 glomol Molaity a. Moles of solute (molt ) solu